1. Successful Services Affiliates (SSA) is considering three new projects, each requiring an initial equipment investment of $21,000. Each project will last for 3 years and produce the following cash flows:
Year Project 1 Project 2 Project 3
1 $ 7,000 $ 9,500 $13,000
2 9,000 9,500 10,000
3 15,000 9,500 11,000
Total $31,000 $28,500 $34,000
The equipment has no salvage value and SSA uses straight-line depreciation. SSA will not accept any project with a payback period over 2 years. Their minimum required rate of return is 12%.
a) Compute each project's payback period, including the most desirable project and the least desirable project using this method. (Round to two decimals).

b) Compute the NPV for each project. (Round to the nearest dollar.)
What is your conclusion?
1
Expert's answer
2012-08-14T09:22:46-0400
& Project 1:
Payback period = 2 +($21,000-$7,000-$9,000)/$15,000=2.3 year
NPV = -$21,000 + $31,000= $10,000

Project 2
Payback period = 2 +($21,000-$9,500-$9,500)/$9,500=2.2 year
NPV = -$21,000 + $28,500= $7,500

Project 3
Payback period = 1 +($21,000-$13,000)/$10,000=1.8year
NPV = -$21,000 + $34,000= $13,000

Project 3 is better for all indices.